# Web 3.0 and Quantum Security: Long-Distance Free-Space QSDC for Global Web 3.0 Networksthanks: These authors contributed equally: Yew Kee Wong, *Yifan Zhou, Xinlin Zhou, Yan Shing Liang, and Zi Yan Li

### v2

@@ I Introduction
Communication Distance
Long-distance (intercontinental)
Moderate distance
Moderate distance
Short to moderate distance
Security Level
High (no key exchange required)
High (key exchange involved)
High (no key exchange involved)
Very high (key exchange fundamental)
Implementation Complexity
Moderate (enhanced by PAT technologies)
High (requires quantum memory)
Moderate (no quantum memory required)
High (requires sophisticated setup)
Suitability for Globalized Web 3.0
Highly suitable
Moderately suitable
Moderately suitable
Less suitable for global scale
Atmospheric Disturbances Resistance
Strong (mitigated by adaptive optics)
Moderate (susceptible to some atmospheric effects)
Moderate (susceptible to some atmospheric effects)
Weak (highly susceptible to atmospheric effects)
@@
Nevertheless, the practical application of quantum communication, particularly in long-distance free-space contexts, faces significant hurdles that must be overcome to fully realize the globalization of Web 3.0 networks. Current quantum communication techniques such as QKD are unable to achieve secure satellite-based long-distance direct data transmission [11, 12, 13, 14]. This research proposes a novel protocol, LF QSDC, aimed at addressing these challenges. LF QSDC is based on memory-free DL04 protocol [15] and incorporates our novel lossless transmission system which includes a Quantum-Aware LDPC coding scheme, PAT technologies, and AQCA.
The memory-free DL04 protocol which FL QSDC employs reduces the complexity of quantum state storage and management, making the system more practical and robust for real-world applications [15] as shown in Table I. Our novel 1uantum-aware LDPC coding scheme, PAT technologies, and AQCA mitigate issues related to atmospheric turbulence and alignment errors, which are prominent in satellite-based free-space quantum communication [16, 17, 18, 19].
@@ II Long-Distance Free-Space QSDC Overview
This section introduces long-distance free-space quan- tum secure direct communication (QSDC). Our investi- gation delves into how advanced technologies can sur- mount existing constraints, rendering QSDC viable for widespread application both in the open air and across space through satellite communication.
@@
II-A State Preparation
The sender, Alice, prepares a sequence of single photons in one of the four time-bin states:
$\begin{split}|T_{0}\rangle=|0\rangle,\quad|T_{1}\rangle&=|1\rangle,\quad|T_{2}\rangle=\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle),\\ \quad|T_{3}\rangle&=\frac{1}{\sqrt{2}}(|0\rangle-|1\rangle),\end{split}$ (1)
where $|0\rangle$ and $|1\rangle$ denote the early and late time bins, respectively [7]. Alice randomly chooses two bits, $a_{0}$ and $a_{1}$, to encode each photon. The encoding rule is as follows [20]:
$(a_{0},a_{1})\rightarrow\left\{\begin{array}[]{ll}|T_{0}\rangle,&\text{if }(a_{0},a_{1})=(0,0),\\ |T_{1}\rangle,&\text{if }(a_{0},a_{1})=(0,1),\\ |T_{2}\rangle,&\text{if }(a_{0},a_{1})=(1,0),\\ |T_{3}\rangle,&\text{if }(a_{0},a_{1})=(1,1).\end{array}\right.$ (2)
Alice also randomly chooses another bit, $a_{2}$, to determine the phase of each photon. The phase modulation rule is as follows [11]:
$a_{2}\rightarrow\left\{\begin{array}[]{ll}\phi=0,&\text{if }a_{2}=0,\\ \phi=\pi,&\text{if }a_{2}=1.\end{array}\right.$ (3)
Alice applies a phase modulator (PM) to each photon according to the value of $a_{2}$. The PM shifts the phase of the late time bin by $\phi$, resulting in the following four phase states [11]:
$\begin{split}|P_{0}\rangle=|0\rangle,\quad|P_{1}\rangle&=|1\rangle,\quad|P_{2}\rangle=\frac{1}{\sqrt{2}}(|0\rangle+e^{i\phi}|1\rangle),\\ \quad|P_{3}\rangle&=\frac{1}{\sqrt{2}}(|0\rangle-e^{i\phi}|1\rangle).\end{split}$ (4)
The phase states are related to the time-bin states by the following transformation [13]:
$|P_{i}\rangle=\frac{1}{\sqrt{2}}(|T_{i}\rangle+e^{i\phi}|T_{i\oplus 2}\rangle),$ (5)
where $\oplus$ denotes addition modulo 4. Alice records the values of $a_{0}$, $a_{1}$, and $a_{2}$ for each photon, and sends the photons to the receiver, Bob, through a free-space channel.
II-B State Transmission
Photons traverse the free-space passage, leading to a range of disruptions including atmospheric disturbances, beam drift, scintillation, and ambient noise. Such disruptions impact the consistency and accuracy of quantum states, potentially leading to inaccuracies or losses in transmission. LF QSDC tackles these obstacles by utilizing a quantum-aware LDPC coding system, merging PAT technologies, and integrating an atmospheric quantum correction algorithm:
@@ II-B State Transmission
    Quantum-Aware LDPC: The LDPC coding system plays a pivotal role in correcting errors inherent in quantum communication. This system is crucial for ensuring the integrity of each transaction in the Web 3.0 era[16].
@@
    PAT Technologies: PAT technologies align quantum signals and maintain a stable communication link, ensuring the continuous operation of decentralized Web 3.0 networks [17].
@@
    Atmospheric Quantum Correction Algorithms: These algorithms address disturbances like turbulence to ensure the reliability of long-distance transmissions crucial for global Web 3.0 networks [6].
This combination of schemes and algorithms ensures a secure, reliable, and efficient quantum communication system, adhering to the requirements of the Web 3.0 era.
II-C State Measurement
Bob receives the photons from Alice and measures them using an interferometer and a single-photon detector (SPD). The interferometer consists of a beam splitter (BS), two PMs, and two mirrors (M). The BS splits each photon into two paths, corresponding to the early and late time bins. The PMs shift the phases of the two paths by $\theta$ and $\theta+\pi$, respectively. The mirrors reflect the photons to the BS, where they recombine and interfere. The SPD detects the photons at the output port of the BS.
The measurement outcome depends on the phase state of the photon and the phase difference $\theta$ between the two paths. Then, Bob randomly chooses the value of $\theta$ for each photon and records the detection results. Bob then announces the values of $\theta$ publicly and discards the results that correspond to $\theta=\frac{\pi}{4}$ or $\theta=\frac{3\pi}{4}$, since these values do not provide any information about the phase of the photon.
II-D Information Extraction
Alice and Bob extract the information from the transmitted and measured photons, using the following steps:
  1. 1.
    Alice and Bob perform sifting, where they compare the values of $a_{2}$ and $\theta$, and keep only the results that satisfy $a_{2}\oplus\theta=0$ or $a_{2}\oplus\theta=1$. This ensures that Bob’s measurement basis is aligned with Alice’s encoding basis and that the phase information is preserved [7].
  2. 2.
    Alice and Bob perform error correction, where they apply the error-correcting codes to correct the bit errors and phase errors in the sifted data. They also perform privacy amplification, applying a hash function to reduce the information leakage to Eve [22].
  3. 3.
    Alice and Bob perform information reconciliation, where they compare the values of $a_{0}$ and $a_{1}$, and extract the common bits as the final secret message. They also perform authentication, where they verify the integrity of the message using a secret key shared beforehand [23][24].

### v3

@@ I Introduction
Communication Distance Long-distance (intercontinental) Moderate distance Moderate distance Short to moderate distance
Security Level High (no key exchange required) High (can transmit with or without key exchange) High (can transmit with or without key exchange) High (key exchange fundamental)
Implementation Complexity Moderate (enhanced by PAT technologies) High (requires quantum memory) Moderate (no quantum memory required) Low (proven by ample experimentation)
Suitability for Globalized Web 3.0 Highly suitable Moderately suitable Moderately suitable Less suitable for global scale
Atmospheric Disturbances Resistance Strong (mitigated by adaptive optics) Moderate (susceptible to some atmospheric effects) Moderate (susceptible to some atmospheric effects) Weak (highly susceptible to atmospheric effects)
@@
Nevertheless, the practical application of quantum communication, particularly in long-distance free-space contexts, faces significant hurdles that must be overcome to fully realize the globalization of Web 3.0 networks. Current quantum communication techniques such as QKD are unable to achieve practical secure satellite-based long-distance direct data transmission yet.
In the evolution of satellite-based QKD, significant strides have been achieved since the launch of the Micius satellite in 2017, a pioneering endeavor that demonstrated the transmission of secure quantum keys over intercontinental distances [3].
Despite these advancements, satellite-based QKD faces persistent challenges, including substantial signal loss, atmospheric interference, and the need for precise alignment and tracking of satellite signals[4, 5, 6].
To address the challenge of a practical, secure, and long-distance quantum communication protocol, this research proposes a novel protocol, LF QSDC. LF QSDC is based on memory-free DL04 protocol [7] and incorporates our novel lossless transmission system which includes a Quantum-Aware LDPC coding scheme, PAT technologies, and AQCA.
Recent advancements in QSDC have shown significant theoretical and experimental progress, enhancing the feasibility of direct quantum communications[8, 9, 10, 11, 12, 13]. Noteworthy developments include the device-independent approach against collective attacks, which has become pivotal for enhancing communication security [14]. Furthermore, the realization of QSDC over 100 km using time-bin and phase quantum states underlines the scalability of these technologies[15].
The DL04 protocol contributes to these advancements by facilitating secure quantum communication without the need for a shared secret key and bypassing the necessity for quantum memory. This protocol ensures the integrity of quantum communication through immediate state preparation and encoding, followed by direct transmission to the receiver. This process is detailed as follows[7]:
State Preparation and Encoding: Similar to other QSDC protocols, the memory-free DL04 requires the sender, Alice, to prepare and encode quantum states. However, instead of storing these states, they are immediately transmitted once prepared.
Quantum State Transmission: Upon preparation, the quantum states are sent directly to the receiver, Bob, who must immediately measure them using predetermined settings agreed upon in advance. This immediate measurement is crucial as it does not allow for the storage and later retrieval of quantum information, thus eliminating the need for quantum memory.
Eavesdropping Detection: The protocol incorporates mechanisms for detecting eavesdropping through quantum state disturbance. Since measurements are performed immediately after transmission, any disturbance caused by an eavesdropper would be detected as anomalies in the measurement results. This approach leverages the fundamental principle of quantum mechanics—any attempt to measure a quantum state will invariably alter it.
Direct Information Transfer: Following successful transmission and measurement, the information is decoded directly from the quantum states. This step confirms the secure transfer of information without the need for intermediate storage, thus negating the requirement for quantum memory.
The memory-free DL04 protocol that FL QSDC employs reduces the complexity of quantum state storage and management, making the system more practical and robust for real-world applications [7] as shown in Table I. Our novel 1uantum-aware LDPC coding scheme, PAT technologies, and AQCA mitigate issues related to atmospheric turbulence and alignment errors, which are prominent in satellite-based free-space quantum communication [16, 17, 18, 19].
@@ II Long-Distance Free-Space QSDC Overview
This section introduces long-distance free-space quantum secure direct communication (QSDC). Our investigation delves into how advanced technologies can sur- mount existing constraints, rendering QSDC viable for widespread application both in the open air and across space through satellite communication.
@@
II-A Quantum State Preparation and Encoding
The LF QSDC protocol starts with Alice preparing and encoding quantum states in a highly structured manner. She uses two parameters, $a_{0}$ and $a_{1}$, to determine the quantum state based on the required information:
Basis Determination ($a_{0}$): - If $a_{0}=0$, the basis is computational, with states $\{|0\rangle,|1\rangle\}$. - If $a_{0}=1$, the basis is superpositional, where $a_{1}$ further dictates the specific state: - $a_{1}=0$: State becomes $\frac{|0\rangle+|1\rangle}{\sqrt{2}}$ $a_{1}=1$: State becomes $\frac{|0\rangle-|1\rangle}{\sqrt{2}}$
Alice also selects a third parameter, $a_{2}$, which applies a phase shift to the encoded states:
- If $a_{2}=0$, the phase states remain as initially encoded. - If $a_{2}=1$, the phase states are swapped: - $|0\rangle+|1\rangle$ becomes $|0\rangle-|1\rangle$ and vice versa.
These encoded quantum states are then transmitted to Bob, who has a crucial role in their measurement.
II-B Measurement Protocols and Two-way Communication
Upon receiving the quantum states, Bob’s task is to measure them accurately, which requires selecting the appropriate basis for measurement based on preliminary information shared by Alice. Bob randomly selects a measurement basis, theta ($\theta$), which dictates his measurement strategy:
- If $\theta=0$, he measures in the basis:
$\{|0\rangle,|1\rangle\}$ (1)
- If $\theta=1$, he measures in the basis:
$\left\{\frac{|0\rangle+|1\rangle}{\sqrt{2}},\frac{|0\rangle-|1\rangle}{\sqrt{2}}\right\}$ (2)
After measurement, Bob announces his choice of $\theta$ and the results of his measurements. Alice and Bob then jointly analyze the outcomes to check if $\theta$ aligns with $a_{2}$ ($a_{2}\oplus\theta=0$ or $1$), ensuring the correct basis was used and the transmission was secure. This process embodies the two-way communication required by the DL04 protocol, where feedback from Bob influences potential retransmissions by Alice or adjustments in the encoding strategy to correct any discrepancies or to enhance security.
II-C Integration of Advanced Quantum Technologies
The effective transmission of quantum states over long distances is fraught with potential disturbances such as atmospheric turbulence, scattering, and absorption. To mitigate these effects and enhance the fidelity of the quantum channel, the following technologies are integrated into our LF QSDC protocol:
@@ II-C Integration of Advanced Quantum Technologies
    Quantum-Aware LDPC Coding: Specifically designed for quantum information, these codes correct errors that occur during the quantum state transmission, thus ensuring that the integrity and secrecy of the data are maintained even over long distances.
@@
    Pointing, Acquisition, and Tracking (PAT) Systems: These technologies are critical for maintaining the alignment of the quantum communication link, especially in dynamic environments such as satellite communications, where precision pointing is crucial for successful data transmission.
@@
    Atmospheric Quantum Correction Algorithms: These algorithms are designed to compensate for the quantum signal degradation caused by the atmosphere. By correcting errors induced by atmospheric turbulence, these algorithms significantly improve the reliability and stability of the quantum channel.